# TabReD: A Benchmark of Tabular Machine Learning in-the-Wild

### v1

@@ Abstract
Benchmarks that closely reflect downstream application scenarios are essential for the streamlined adoption of new research in tabular machine learning (ML). In this work, we examine existing tabular benchmarks and find two common characteristics of industry-grade tabular data that are underrepresented in the datasets available to the academic community. First, tabular data often changes over time in real-world deployment scenarios. This impacts model performance and requires time-based train and test splits for correct model evaluation. Yet, existing academic tabular datasets often lack timestamp metadata to enable such evaluation. Second, a considerable portion of datasets in production settings stem from extensive data acquisition and feature engineering pipelines. For each specific dataset, this can have a different impact on the absolute and relative number of predictive, uninformative, and correlated features, which in turn can affect model selection. To fill the aforementioned gaps in academic benchmarks, we introduce TabReD – a collection of eight industry-grade tabular datasets covering a wide range of domains from finance to food delivery services. We assess a large number of tabular ML models in the feature-rich, temporally-evolving data setting facilitated by TabReD. We demonstrate that evaluation on time-based data splits leads to different methods ranking, compared to evaluation on random splits more common in academic benchmarks. Furthermore, on the TabReD datasets, MLP-like architectures and GBDT show the best results, while more sophisticated DL models are yet to prove their effectiveness.
@@ 1 Introduction
During several recent years, research on tabular machine learning has grown rapidly. Plenty of works have proposed neural network architectures [33, 52, 19, 20, 21, 10, 11] that are competitive or even superior to “shallow” GBDT models [41, 30, 13], which has strengthened research interest in the field. Furthermore, specialized workshops devoted to table representation learning are organized on the top-tier ML conferences [^1] Table Representation Learning Workshop, NeurIPS, which highlights the importance of this research line to the community.
Rapid progress on a particular problem requires a comprehensive set of benchmarks. It is important that benchmarks reflect the characteristics and specifics of downstream applications to ensure seamless adoption of research progress [43, 51, 34]. In this work we rigorously investigate tabular datasets from academic benchmarks and find that two data characteristics common in industrial tabular ML applications are underrepresented in the current academic benchmarks.
@@
First, in practice, datasets are often split into train/validation/test parts according to timestamps of datapoints [23, 47, 8, 27, 25] to account for potential distribution shifts that can naturally occur in real applications. In fact, even among academic benchmarks, there is a certain number of datasets with strong time dependencies between instances (e.g. electricity market prediction [50], flight delay estimation [7], bike sharing demand [49], and others). However, even in such cases, random splits are used in papers instead of time-based splits. Moreover, timestamp or other task-appropriate split metadata is often not available.
@@
We also find that datasets with large numbers of predictive features and extensive feature engineering are scarce in academic benchmarks. Conversely, such feature-rich datasets are common in many industrial settings [17, 45, 28, 5, 52], but they are often proprietary and unavailable to the academic community.
As a bonus point, we also identify eleven data-leakage issues and the use of non-tabular, synthetic and untraceable datasets in the academic benchmarks. All TabReD datasets are free from such issues.
@@
To fill the gap in existing academic benchmarks and represent in-the-wild industrial tabular ML applications better, we introduce the TabReD benchmark – a collection of eight datasets, all drawn from real-world industrial applications with tabular data. All TabReD datasets come with time-based splits into train, validation and test parts. For a given application, it means that test datapoints appeared later in time than training datapoints. This approach takes into account the gradual “temporal shift” between training and test objects (i.e. the shift between “older” and “newer” objects), and we confirm the importance of such time-based splits in experiments. Furthermore, because of additional investments in data acquisition and feature engineering, all datasets in TabReD have more features. This stems from adopting the preprocessing steps from production ML pipelines and Kaggle competition forums, where extensive data engineering is often highly prioritized.
We evaluate numerous tabular models on the TabReD benchmark. We find that the general trend of tabular DL progress holds there, but the improvements brought by model architecture advances are less pronounced and simpler MLP-like and GBDT solutions prove to be more universally practical and performant on TabReD.
@@
  • We introduce TabReD – a collection of eight industry-grade tabular datasets that span a wide range of domains from finance to food delivery services (three classification and five regression problems). The datasets in TabReD exhibit two important practical properties that, as we show in our analysis, are underrepresented in existing academic benchmarks. First, the TabReD datasets are split into train/validation/test sets according to timestamps of objects, which we show to be crucial for correct evaluation. Second, the TabReD datasets have more features as a result of additional effort put into data acquisition and feature engineering.
  • We evaluate a large number of tabular models on TabReD. We find that, in the feature-rich, time-evolving setting facilitated by TabReD: (1) GBDT and MLPs with embeddings [20] demonstrate the best average performance; (2) more complex DL models turn out to be less effective; (3) the difference in metrics between various methods is less pronounced than on academic benchmarks used in prior work.
  • We demonstrate that correct evaluation on datasets with temporally shifted validation and test sets is crucial as it leads to significant differences in rankings and relative performance of methods, compared to commonly used random-split-based evaluation. In particular, we observe that XGBoost performance margin diminishes in correct evaluation setups.
@@ 2 Related Work
Tabular deep learning is a dynamically developing research area with the recent works proposing new model architectures [33, 52, 19, 20, 21, 10, 11], regularizations [26], pretraining pipelines [42, 6] and other solutions [24]. Since the common way to justify the usage of new approaches is to empirically compare them against the baselines, the choice of the benchmarks for evaluation is critical.
@@
Tabular Data Benchmarks. Since tabular tasks occur in a large number of applications from various domains, there is no single dataset that would be sufficient for evaluation. A typical tabular DL paper reports the results on several tasks from different domains, usually coming from one of the public data repositories. The two traditional data sources for ML problems are the UCI  [^2] https://archive.ics.uci.edu and OpenML [^3] https://www.openml.org repositories, currently holding thousands of datasets. Unfortunately, datasets available in public repositories do not cover all tabular ML use cases. In particular, we find that industry-grade tabular ML datasets are underrepresented in these public repositories.
@@
Another source of datasets is the Kaggle [^4] https://www.kaggle.com platform, which hosts a plethora of ML competitions, including ones with tabular data. Datasets from competitions are often provided by people solving particular real-world problems, making Kaggle an attractive source of datasets for tabular ML research. Surprisingly, many benchmarks rely on UCI and OpenML and underutilize tabular datasets from Kaggle. For example, out of 100 academic datasets that we analyze in subsection 3.1, only four come from Kaggle competitions. To construct TabReD, we source datasets from Kaggle tabular data competitions, and we also introduce four new datasets from real ML production systems at a large tech company.
@@
Tabzilla [39] and the Grinsztajn et al. [22] benchmark have gained adoption in the research community. For example, Gorishniy et al. [21] and Chen et al. [12] demonstrate models competing with GBDTs on the Grinsztajn et al. [22] benchmark, Feuer et al. [16] use a mix of Tabzilla and Grinsztajn et al. [22] to test an approach for scaling-up TabPFN [24]. Both benchmarks primarily rely on the OpenML repository as a source of datasets and filter datasets semi-automatically based on metadata like size and baseline performance. In our work, we look closer at all the datasets and find data-leakage issues and non-tabular, synthetic and anonymous/unknown datasets sometimes “sneaking” through automatic filters. Furthermore, many datasets in these two benchmarks lack task-appropriate test splits and use random splits for measuring the performance. We show that ignoring temporal data drift in evaluation could lead to significant differences in methods ranking and relative performance.
@@
TableShift [18] and WildTab [35] propose tabular benchmarks with distribution shifts between train/test subsets. These benchmarks are closer in spirit to TabReD, as both mention the importance of moving towards more realistic evaluation scenarios. However, both benchmarks focus on out-of-distribution robustness and domain generalization methods comparison, not including many tabular-data specific methods. We study a broader set of methods, including recent SoTA tabular neural networks. Furthermore, both benchmarks consider more “extreme” shifts, compared to more ubiquitous gradual temporal shift which is present in all TabReD datasets.
@@
Benchmarking Under Temporal Shift. Wild-Time [53] proposed a benchmark consisting of five datasets with temporal shift and identified a 20% performance drop due to temporal shift on average. However, tabular data was not the focus of the Wild-Time benchmark.
@@
The presence of temporal shift and the importance of evaluating models under temporal shifts was discussed in many research and application domains including approximate neighbors search [8], recommender systems [44], finance applications [47, 23], health insurance [27] and in general ML systems best practices [25].
3 TabReD Benchmark
In this section, we take a closer look at existing benchmarks for tabular deep learning and summarize the characteristics that make them different from datasets in industrial ML applications. Then we introduce the TabReD benchmark that aims to address these differences.
3.1 A Closer Look at Existing Tabular ML Benchmarks
@@
We collect unique regression and classification datasets from academic tabular ML benchmarks [20, 22, 21, 39, 12, 18], this results in a set of 100 datasets. We carefully assess each dataset and summarize our findings in Table 1. We also provide detailed meta-data collected in the process with short descriptions of tasks, original data sources, data quality issues and notes on temporal splits in the Appendix E and as a separate table in the supplementary material. Our main findings are summarized below.
@@
Table 1: The landscape of existing tabular machine learning benchmarks compared to TabReD. We report median dataset sizes, number of features, the number of datasets with various issues. The “Time-splits” column is reported only for the datasets without issues. We see that the datasets semi-automatically gathered from OpenML (Tabzilla and Grinsztajn et al. [22]) contain more quality issues. Furthermore, no benchmark besides TabReD focuses on temporal-shift based evaluation, less than half of datasets in each benchmark have timestamp metadata needed for time-based validation availability.
@@
   * – the original dataset, introduced in [38] has the canonical OOD split, but the standard IID split commonly used contains time-based leakage.
@@
Grinsztajn et al. [22] 16,679 13 7 / 44 1 / 44 7 / 44 22 5
Tabzilla [39] 3,087 23 3 / 36 6 / 36 12 / 36 12 0
WildTab [35] 546,543 10 1* / 3 1 / 3 0 / 3 1 1
TableShift [18] 840,582 23 0 / 15 0 / 15 0 / 15 15 8
Gorishniy et al. [21] 57,909 20 1* / 10 1 / 10 0 / 10 7 1
@@
Data Leakage, Synthetic and Non-Tabular Datasets. First, we find that a considerable number of tabular datasets have some form of data leakage (11 out of 100). Leakage stems from data preparation errors, near-duplicate instances or inappropriate data splits used for testing. A few of these leakage issues have been reported in prior literature [21], but as there are no common protocols for deprecating datasets in ML [37], datasets with leakage issues are still used. For some datasets the data source is untraceable, or the data is known to be synthetic without the generation process details or description – there are 13 such datasets. Last, we find that 25 datasets used in academic benchmarks are not inherently tabular by the categorization proposed in Kohli et al. [34]. These datasets either represent raw data stored in a table form (e.g. flattened images) or homogenous features extracted from some raw data source.
@@
Dataset Size and Feature Engineering. We find that most datasets from the academic benchmarks have less than 60 features and less than a few hundred instances available. Many academic datasets come from publicly available data, which often contain only high-level statistics (e.g. only the source and destination airport and airline IDs for the task of predicting flight delays in the dataset by Ballesteros [7]). In contrast, many in-the-wild industrial ML applications utilize as much information and data as possible [e.g. 17, 45, 28]. Unfortunately, not many such datasets from in-the-wild applications are openly available for research. Kaggle competitions come closest to this kind of industry-grade tabular data, but using competition data is less common in current academic benchmarks (only 4 datasets are from Kaggle competitions).
@@
Lack of canonical splits or timestamp metadata. All benchmarks except the ones focused on distribution shift do not discuss the question of data splits used for model evaluation, beyond standard experimental evaluation setups (e.g. random split proportions or cross-validation folds). We find that 53 existing datasets (excluding datasets with issues) potentially contain data drifts related to the passage of time, as the data was collected over time. It is a standard industry practice to use time-based splits for validation in such cases [44, 47, 27, 25]. However, only 15 datasets have timestamps available for such splits.
3.2 Constructing the TabReD Benchmark

### v4

# TabReD: Analyzing Pitfalls and Filling the Gaps in Tabular Deep Learning Benchmarks

@@ Abstract
Advances in machine learning research drive progress in real-world applications. To ensure this progress, it is important to understand the potential pitfalls on the way from a novel method’s success on academic benchmarks to its practical deployment. In this work, we analyze existing tabular benchmarks and find two common characteristics of tabular data in typical industrial applications that are underrepresented in the datasets usually used for evaluation in the literature. First, in real-world deployment scenarios, distribution of data often changes over time. To account for this distribution drift, time-based train/test splits should be used in evaluation. However, popular tabular datasets often lack timestamp metadata to enable such evaluation. Second, a considerable portion of datasets in production settings stem from extensive data acquisition and feature engineering pipelines. This can have an impact on the absolute and relative number of predictive, uninformative, and correlated features compared to academic datasets. In this work, we aim to understand how recent research advances in tabular deep learning transfer to these underrepresented conditions. To this end, we introduce TabReD – a collection of eight industry-grade tabular datasets. We reassess a large number of tabular ML models and techniques on TabReD. We demonstrate that evaluation on time-based data splits leads to different methods ranking, compared to evaluation on random splits, which are common in current benchmarks. Furthermore, simple MLP-like architectures and GBDT show the best results on the TabReD datasets, while other methods are less effective in the new setting.
@@ 1 Introduction
During several recent years, research on tabular machine learning has grown rapidly. Plenty of works have proposed neural network architectures (Klambauer et al., 2017; Wang et al., 2020; Gorishniy et al., 2021; 2022; 2024; Chen et al., 2023a; b; Ye et al., 2024) that are competitive or even superior to “shallow” GBDT models (Prokhorenkova et al., 2018; Ke et al., 2017; Chen & Guestrin, 2016), which has strengthened research interest in the field. Furthermore, specialized workshops devoted to table representation learning are organized on the top-tier ML conferences [^1] Table Representation Learning Workshop, NeurIPS, which highlights the importance of this research line to the community.
An important benefit of machine learning research is the practical application of novel methods developed in academia. However, some conditions encountered in real-world deployment can challenge the benefits of the proposed methods. To study how different methods perform under these conditions, one needs a representative group of datasets. In this work, we take a closer look at existing tabular benchmarks. We investigate each dataset and identify some of them as synthetic, untraceable or non-tabular. Moreover, we find and highlight eleven datasets containing leaks. But most importantly, in our analysis, we find two data characteristics common in industrial tabular ML applications that are underrepresented in the current benchmarks.
@@
First, in practice, data is subject to a gradual temporal shift. To account for this, in practice, datasets are often split into train/validation/test parts according to timestamps of datapoints (Herman et al., 2024; Stein, 2002; Baranchuk et al., 2023; Ji et al., 2023; Huyen, 2022). This scenario differs from the types of distribution shift covered in (Gardner et al., 2023), because the temporal shift is gradual and the time data representing the degree of this shift is available to the developer. In fact, even among currently used benchmarks, there is a certain number of datasets with strong time dependencies between instances (e.g. electricity market prediction (van Rijn, 2020), flight delay estimation (Ballesteros, 2019), bike sharing demand (van Rijn, 2014), and others). However, even in such cases, random splits are used in papers instead of time-based splits. Moreover, timestamp or other task-appropriate split metadata is often simply not available.
@@
Second, we find that datasets with large numbers of predictive features and extensive feature engineering are scarce in tabular benchmarks. Conversely, such feature-rich datasets are common in many industrial settings (Fu & Soman, 2021; Simha, 2020; Kakade, 2021; Anil et al., 2022; Wang et al., 2020), but they are often proprietary and unavailable to the academic community.
@@
In light of these discoveries, we study a question of transferability of novel methods in tabular DL to these underrepresented conditions. To answer this question and fill the gap in existing benchmarks, we introduce the TabReD benchmark – a collection of eight datasets, all drawn from real-world industrial applications with tabular data. All TabReD datasets come with time-based splits into train, validation and test parts. Furthermore, because of additional investments in data acquisition and feature engineering, all datasets in TabReD have more features on average than in previous benchmarks. This stems from adopting the preprocessing steps from production ML pipelines and Kaggle competition forums, where extensive data engineering is often highly prioritized.
We evaluate numerous tabular methods on the TabReD benchmark. We find that most of the tested novel architectures and techniques that show improvements on previously available benchmarks do not show benefits on our datasets, while simple MLP architecture with embeddings and GBDT methods show top performance across the new benchmark.
@@
  • We analyze the existing tabular benchmarks in academia and find that temporally-evolving and feature-rich datasets are underrepresented in commonly used benchmarks. We also find data leakage issues, use of non-tabular, synthetic or untraceable datasets.
  • We introduce TabReD – a collection of eight industry-grade tabular datasets that span a wide range of domains, from finance to food delivery services. With TabReD we increase the coverage of industrial tabular ML use-cases in academic benchmarks. All TabReD datasets are a result of industrial feature engineering pipelines and use time-based evaluation.
  • We evaluate a large number of tabular DL techniques on TabReD. We find that, in the feature-rich, time-evolving setting facilitated by TabReD: (1) GBDT and MLPs with embeddings (Gorishniy et al., 2022) demonstrate the best average performance; (2) more complex DL methods turn out to be less effective. We demonstrate that correct evaluation on datasets with temporally shifted validation and test sets is crucial as it leads to significant differences in rankings and relative performance of methods, compared to commonly used random-split based evaluation. In particular, we observe that XGBoost performance margin diminishes in correct evaluation setups.
@@ 2 Related work
Tabular deep learning is a dynamically developing research area, with the recent works proposing new model architectures (Klambauer et al., 2017; Wang et al., 2020; Gorishniy et al., 2021; 2022; 2024; Chen et al., 2023a; b), regularizations (Jeffares et al., 2023), (pre-)training pipelines (Bahri et al., 2021; Rubachev et al., 2022; Lee et al., 2024) and other solutions (Hollmann et al., 2023). Since the common way to justify the usage of new approaches is to empirically compare them against the baselines, the choice of the benchmarks for evaluation is critical.
@@
Tabular Data Benchmarks. Since tabular tasks occur in a large number of applications from various domains, there is no single dataset that would be sufficient for evaluation. A typical tabular DL paper reports the results on several tasks from different domains, usually coming from one of the public data repositories. The two traditional data sources for ML problems are the UCI  [^2] https://archive.ics.uci.edu and OpenML [^3] https://www.openml.org repositories, currently holding thousands of datasets. In particular, we find that some conditions of industrial tabular ML applications are underrepresented in these public repositories.
@@
Another source of datasets is the Kaggle [^4] https://www.kaggle.com platform, which hosts a plethora of ML competitions, including ones with tabular data. Datasets from competitions are often provided by people solving particular real-world problems, making Kaggle an attractive source of datasets for tabular ML research. Nevertheless, many popular benchmarks rely on UCI and OpenML and underutilize tabular datasets from Kaggle. For example, out of 100 commonly used datasets that we analyze in section 3, only four come from Kaggle competitions. To construct TabReD, we utilize several datasets from Kaggle tabular data competitions, and we also introduce four new datasets from real ML production systems at a large tech company.
@@
Tabzilla (McElfresh et al., 2023) and the Grinsztajn et al. (2022) benchmark have gained adoption in the research community. For example, such papers as Gorishniy et al. (2024), Chen et al. (2023b), Feuer et al. (2024), demonstrate their performance on these benchmarks. Both benchmarks primarily rely on the OpenML repository as a source of datasets, and filter datasets semi-automatically based on metadata like size and baseline performance. In our work, we look closer at all the datasets and find data-leakage issues and non-tabular, synthetic and anonymous/unknown datasets sometimes “sneaking” through automatic filters. Furthermore, these datasets do not represent conditions of temporal shift and extensive feature engineering that are common in practical applications.
@@
TableShift (Gardner et al., 2023) and WildTab (Kolesnikov, 2023) propose tabular benchmarks with distribution shifts between train/test subsets. However, both benchmarks focus on out-of-distribution robustness and domain generalization methods comparison, not how many tabular-data specific methods perform in the new setting. We study a broader set of methods, including recent SoTA tabular neural networks. Furthermore, both benchmarks consider more “extreme” shifts, compared to the more ubiquitous gradual temporal shift which is present in all TabReD datasets.
@@
Benchmarking Under Temporal Shift. Wild-Time (Yao et al., 2022) proposed a benchmark consisting of five datasets with temporal shift and identified a 20% performance drop due to temporal shift on average. However, tabular data was not the focus of the Wild-Time benchmark.
@@
The presence of temporal shift and the importance of evaluating models under temporal shifts was discussed in many research and application domains including approximate neighbors search (Baranchuk et al., 2023), recommender systems (Shani & Gunawardana, 2011), finance applications (Stein, 2002; Herman et al., 2024), health insurance (Ji et al., 2023) and in general ML systems best practices (Huyen, 2022).
3 A closer look at the existing tabular ml benchmarks
In this section, we take a closer look at the existing benchmarks for tabular deep learning. We analyze dataset sizes, number of features, the presence of temporal shift and its treatment. We also point out the issues of some datasets that we find notable. The first such issue is the presence of leakage. We believe that an inadvertent usage of datasets with leaks has a negative impact on the quality of evaluation. We provide a list of such datasets in the corresponding section below. The second issue is the untraceable or synthetic nature of the data. Without knowing the source of the data and what it represents, it is unclear how transferable are advances on these datasets. The third issue is the usage of the data that belongs to other domains, e.g. image data flattened into an array of values. While such datasets correspond to a valid and useful task, it is unclear how useful are advances on such datasets in practice, since other domain-specific methods usually perform significantly better for this type of data.
@@ 3 A closer look at the existing tabular ml benchmarks
Table 1 summarizes our analysis of 100 unique classification and regression datasets from previously available benchmarks (Gorishniy et al., 2022; Grinsztajn et al., 2022; Gorishniy et al., 2024; McElfresh et al., 2023; Chen et al., 2023b; Gardner et al., 2023). We also provide detailed meta-data collected in the process with short descriptions of tasks, original data sources, data quality issues and notes on temporal splits in the Appendix E. Our main findings are as follows.
@@
Table 1: The landscape of existing tabular machine learning benchmarks compared to TabReD. We report median dataset sizes, number of features, the number of datasets with various issues. The “Time-splits” column is reported only for the datasets without issues. We see that the datasets semi-automatically gathered from OpenML (Tabzilla and Grinsztajn et al. (2022)) contain more quality issues. Furthermore, no benchmark besides TabReD focuses on temporal-shift based evaluation and less than half of datasets in each benchmark have timestamp metadata needed for time-based validation availability.
@@
   * – the original dataset, introduced in (Malinin et al., 2021) has the canonical OOD split, but the standard IID split commonly used contains time-based leakage.
@@
Grinsztajn et al. (2022) 16,679 13 7 / 44 1 / 44 7 / 44 22 5
Tabzilla (McElfresh et al., 2023) 3,087 23 3 / 36 6 / 36 12 / 36 12 0
WildTab (Kolesnikov, 2023) 546,543 10 1* / 3 1 / 3 0 / 3 1 1
TableShift (Gardner et al., 2023) 840,582 23 0 / 15 0 / 15 0 / 15 15 8
Gorishniy et al. (2024) 57,909 20 1* / 10 1 / 10 0 / 10 7 1
@@
Data Leakage, Synthetic and Non-Tabular Datasets. First, a considerable number of tabular datasets have some form of data leakage (11 out of 100). Leakage stems from data preparation errors, near-duplicate instances or inappropriate data splits used for testing. A few of these leakage issues have been reported in prior literature (Gorishniy et al., 2024), but as there are no common protocols for deprecating datasets in ML (Luccioni et al., 2022), datasets with leakage issues are still used. Here is the list of datasets where we identified leaks: eye_movements, visualizing_soil, Gesture Phase, sulfur, artificial-characters, compass, Bike_Sharing_Demand, electricity, Facebook Comments Volume, SGEMM_GPU_kernel_performance, Shifts Weather (in-domain-subset). For some datasets the data source is untraceable, or the data is known to be synthetic without the generation process details or description – there are 13 such datasets. Last, we find that 25 datasets used in current benchmarks are not inherently tabular by the categorization proposed in Kohli et al. (2024). These datasets either represent raw data stored in a table form (e.g. flattened images) or homogenous features extracted from some raw data source.
@@
Dataset Size and Feature Engineering. We find that most datasets from the current benchmarks have less than 60 features and less than a hundred thousand instances available. Many datasets used in academia come from publicly available data, which often contain only high-level statistics (e.g. only the source and destination airport and airline IDs for the task of predicting flight delays in the dataset by Ballesteros (2019)). In contrast, many in-the-wild industrial ML applications utilize as much information and data as possible (e.g. Fu & Soman, 2021; Simha, 2020; Kakade, 2021). Unfortunately, not many such datasets from in-the-wild applications are openly available for research. Kaggle competitions come closest to this kind of industry-grade tabular data, but using competition data is less common in current benchmarks (only 4 datasets are from Kaggle competitions). To highlight the difference of previous benchmarks with ours, we provide information on correlated features in subsection A.2.
@@
Lack of canonical splits or timestamp metadata. All benchmarks except the ones focused on distribution shift do not discuss the question of data splits used for model evaluation, beyond standard experimental evaluation setups (e.g. random split proportions or cross-validation folds). We find that 53 existing datasets (excluding datasets with issues) potentially contain data drifts related to the passage of time, as the data was collected over time. It is a standard industry practice to use time-based splits for validation in such cases (Shani & Gunawardana, 2011; Stein, 2002; Ji et al., 2023; Huyen, 2022). However, only 15 datasets have timestamps available for such splits.
4 Constructing the TabReD benchmark